BROADBAND VELOCITY MODULATION SPECTROSCOPY OF MOLECULAR IONS FOR USE IN THE JILA ELECTRON EDM EXPERIMENT

\underline{DANIEL N. GRESH}, KEVIN C. COSSEL, ERIC A. CORNELL, and JUN YE, JILA, National Institute of Standards and Technology and University of Colorado Department of Physics, University of Colorado, Boulder, Colorado 80309-0440, USA.

The JILA electron electric dipole moment (eEDM) experiment will use a low-lying, metastable ^{3}Δ_{1} state in trapped molecular ions of HfF^{+} or ThF^{+}.^{a} Prior to this work, the low-lying states of these molecules had been investigated by PFI-ZEKE spectroscopy.^{b} However, there were no detailed studies of the electronic structure. The recently developed technique of frequency comb velocity modulation spectroscopy (VMS)^{c} provides broad-bandwidth, high-resolution, ionsensitive spectroscopy, allowing the acquisition of 150 cm^{−1} of continuous spectra in 30 minutes over 1500 simultaneous channels. By supplementing this technique with cw-laser VMS, we have investigated the electronic structure of HfF^{+} in the frequency range of 9950 to 14600 cm^{−1}, accurately fitting and assigning 16 rovibronic transitions involving 8 different electronic states including the X^{1}Σ^{+} and a^{3}Δ_{1} states.^{d} In addition, an observed ^{3}Π_{0+} state with coupling to both the X and a states has been used in the actual eEDM experiment to coherently transfer population from the rovibronic ground state of HfF^{+} to the eEDM science state. Furthermore, we report on current efforts of applying frequency comb VMS at 700 - 900 nm to the study of ThF^{+}, which has a lower energy ^{3}Δ_{1} state and a greater effective electric field, and will provide increased sensitivity for a measurement of the eEDM.

aA. E. Leanhardt et. al., Journal of Molecular Spectroscopy 270, 1-25 (2011).

bB. J. Barker, I. O. Antonov, M. C. Heaven, K. A. Peterson, Journal of Chemical Physics 136, 104305 (2012).

cL. C. Sinclair, K. C. Cossel, T. Coffey, J. Ye, E. A. Cornell, Physical Review Letters 107, 093002 (2011).

dK.C. Cossel et. al., Chemical Physics Letters 546, 1-11 (2012).